OpenStudy (anonymous):

can someone check my answer medal included!! A gardener weeds 3/4 of a flower bed in 6 min. How much of the flower bed did the gardener weed in 1 min. my answer) 1/9

OpenStudy (sleepyjess):

if 3/4 takes him 6 min that means 1/4 takes 2 min \(\dfrac{1}{4}/\dfrac{2}{1}=\dfrac{1}{4}*\dfrac{1}{2}=\dfrac{1}{8}\)

OpenStudy (anonymous):

so it 1/8